# Modelling SARS-CoV-2 epidemics via compartmental and cellular automaton SEIRS model with temporal immunity and vaccinationthanks: The manuscript is dedicated to the memory of Prof. Stefan Sokołowski, outstanding scientist, teacher, colleague and the friend

### 23 Mar 2026 v1

@@ 1 Introduction and the model
The minimalistic SEIRS compartmental model, which incorporates the effects of vaccination and of a loss of immunity, is illustrated in figure 1. It contains four compartments, classified by the type of individuals they contain: susceptible to virus, $S$; unidentified infective, $E$; identified isolated infective, $I$; and recovered non-infective, $R$. The same letters, $S$, $E$, II, and $R$, denote respective fractions of individuals they contain.
@@
Infecting (transfer of an individual from $S$ to $E$ compartment) occurs with the rate of $\beta$. To keep the model simple, we do not introduce a separate compartment for exposed individuals [23, 48, 41, 42]. Therefore, the infecting rate, $\beta$, is affected by both social distancing, and by keeping sanitary norms, but these two factors cannot be separated in our model. Infected individual became instantly infective, as follows from the known facts of infectivity during the incubation period [66]. Infective individuals are identified (transferred from $E$ to $I$), via testing, with a rate of $\alpha$. Again, not to overload a model with too many processes, we consider the case of an ideally disciplined society, where the identified infective individual becomes instantly isolated, either in home or in a hospital. The difference in curing in both cases is rather minor, save for fatal and severe forms. Therefore, we do not distinguish such patients, in contrast to some other models [23, 41]. Both unidentified and isolated infective individuals are supposed to recover, acquire temporal immunity, and loose their infectivity (transferred to $R$ compartment) with the same recovery rate of $\gamma$. Vaccination provides a shortcut from the $S$ compartment straight into the $R$ one. We note, that vaccination in our model is explicit, and changed dynamically with the time evolution of the system, contrary to a static vaccination case [10], where the fraction of immune individuals is set constant. We neglect the vaccine activation period, and do not separate such factors as population coverage by vaccination and vaccine efficacy [41], since both are combined into a single vaccination rate $\omega$. In this way, all individuals residing in the $R$ compartment are assumed to be temporarily immunized. Upon loss of immunity these are transferred back into the susceptible group $S$ with the loss of immunity rate $\varphi$. These “returned” individuals may be subject to reinfection [18, 59, 27]. This may occur due to the mutation of the SARS-CoV-2 virus, e.g., into Alpha, Delta, or Omicron variant, but this can be taken into account in our model only implicitly, via the increase of the loss of immunity rate of $\varphi$. The model also neglects the birth and death rates.
@@
The rates $\beta$, $\alpha$, $\gamma$, $\varphi$ and $\omega$ can be related to the real life statistical data. If the population size is $N$ individuals, then the absolute number of susceptible, unidentified infective, identified isolated infective and recovered non-infective individuals are $N_{S}=SN$, $N_{E}=EN$, $N_{I}=IN$ and $N_{R}=RN$, respectively. Assuming a time unit equal to a single day, the absolute number of individuals infected per day is $\beta N_{S}(N_{E}/N)$. The coefficient $\alpha$ reflects how widely the population is covered by appropriate medical testing and defines the number of newly identified infective individuals per day, $\alpha N_{E}$. Such type of data should be available from medical institutions. Recovery rate $\gamma$, is inversely proportional to the average period of infectivity, for which much data have been systemized in reference [77]. There is some distribution of the results depending on many factors, but a reasonable average estimate of $14$ days is often cited [62, 19, 6, 69, 51]. This leads to the following estimate: $\gamma=1/14$. The loss of immunity rate, $\varphi$, inversely proportional to the time interval between (re)infections, is difficult to estimate as an unique average value [74]. Though the early studies reported the time interval for the loss of immunity of 2–3 months [4], 3–4 months [7], as well as long lasting immune memory [21], later works emphasize the role of vaccination, advanced age, the presence of comorbidities [29]. The presence of chronic diseases, namely: ischemic and inflammatory heart disease, dysrhythmias, venous thromboembolism, cerebrovascular diseases are found as an important factor as well [75]. Therefore, one may face the need to fix the time interval to a certain value, e.g., $120$ days, yielding the estimate for $\varphi=1/120$. Finally, the vaccination rate, $\omega$, can be estimated from the absolute number of individuals vaccinated per day, $\omega N_{S}$, the information available from medical institutions.
@@
One can clearly identify the effect of the presence of isolated $I$ and recovered $R$ individuals in this model by introducing cumulative fractions of all uninfected individuals, $S^{\prime}=S+R$, and all infected ones, $E^{\prime}=E+I$, and the normalization is: $S^{\prime}+E^{\prime}=1$. The equations set (1.1)–(1.4) can be rewritten as a single equation for $E^{\prime}$
@@
which has the same form as the one for the $SIS$ epidemiology model [36], when $S$ and $I$ are substituted via $S^{\prime}$ and $E^{\prime}$, respectively, but with the reduced number of transmission acts. The latter are given by the product $(E^{\prime}-I)(S^{\prime}-R)$, where both infective and susceptible parties are lessened by $I$ and $R$, respectively. The equation (1.6) is, obviously, not self-sufficient, as one needs to complement it by the equations providing the time evolution for the $I$ and $R$ fractions. To do so, one may attempt either to rewrite the equations set (1.1)–(1.4) in terms of four variables $S^{\prime}$, $E^{\prime}$, $I$ and $R$, or to use certain approximations to express $E$ and $R$ fractions via $S^{\prime}$ and $I^{\prime}$ variables, similarly as suggested in reference [37] for the SEIRS model without the immunity and vaccination considered there.
@@ 2 Fixed points and their stability
where $R_{0}$ has a meaning of the basic reproductive number. Let us note that in the limit case of $\alpha=\omega=0$ (no identification and no vaccination), the expression $R_{0}=\beta/\gamma$ for the SIS model is retrieved (in this case the $E$ fraction of the current model serves as the $I$ fraction in the SIS model). Taking into account the assumption (2.6), the expression (2.16) indicates the way of bringing the basic reproductive number $R_{0}$ down by means of both decrease of the transmission rate $\beta$ and by an increase of the identification $\alpha$ and vaccination $\omega$ rates. One can rewrite the condition $R_{0}\leqslant 1$ for the DF state in any of the following three forms
@@
at the time instance $t$ depend on all variables, $S$, $E$, II and $R$ at the same instance $t$. The source of infection is the $E$ fraction, because in the current model, depicted in figure 1, we consider an ideal case of instant isolation of identified infective individuals into a group $I$. On the other hand, part of the latter are isolated in the hospitals and at home, therefore $I$ reflects the load put on the health system. Therefore, we concentrate on the time evolutions, $E(t)$ and $I(t)$, of these two fractions. These are examined at various initial values of $E_{0}$ and at various $\beta$, $\alpha$ and $\omega$ rates.
@@ 4 Cellular automaton simulations
To take into account the spatial arrangement of individuals within the SEIRS model as defined in figure 1, we consider their simplest 2D arrangement: a simple square lattice with a neighbourhood size for each individual, given by the number $q$ of its neighbours. The setup is similar to that being discussed for the case of a simpler SIS model in reference [36]. In particular, the vertex set of a graph is $\mathbb{Z}^{2}$, and each individual is characterized by a neighbourhood radius $R_{q}>0$, such that the neighbourhood of a given $k\in\mathbb{Z}^{2}$ is defined as: $k^{\prime}\sim k$ whenever $|k^{\prime}-k|\leqslant R_{q}$. The choice of $R_{q}$ is made to obtain the neighbourhood size equal to the previously chosen neighbourhood size $q$. Each $k$ is associated with an individual, that is characterized by its state $s_{k}$, which takes any value from a set of $\{s,e,i,r\}$, the letters correspond to the susceptible, unidentified infected, identified infected and recovered states, respectively. The evolution of the system of $N$ such individuals is driven by the algorithm, based on transition rules between the states of each individual in the SEIRS model, depicted schematically in figure 1. The principal difference with the compartmental realization of this model is that the transmission occurs locally, within the sphere of the radius $R_{q}$ around each infected individual.
We consider a square lattice of $N=700^{2}=490\,000$ individuals which brings the model to a realistic size of a medium-sized town. The periodic boundary conditions are applied along both Cartesian axes. The set of neighbourhood sizes ranges from $q=N$ (the limit of the compartmental model) down to $q=4$ (von Neumann neighbourhood). In this way the effect of quarantine is considered explicitly, by reducing the neighbourhood size $q$, and can be partly separated from the probability to contract a disease. At each time instance, we made $N$ random choices of individuals, and their states are updated according to their current state and the states of their neighbours. This levels up the time discretization with the $\Delta t$ of the numeric integration (3.1). We employed the asynchronous realization of the cellular automaton algorithm, when the change of $s_{k}$ occurs immediately, and affects the rest of $N$ updating attempts performed at the same time instance.
We focus here on the time evolution of the fraction of identified infected individuals, $I(t)$, as it characterizes the load put on the medical system of an imaginary town. The strength of quarantine measures is defined by restrictions imposed on social contacts defined via the neighbourhood size $q$. It is fixed throughout the system of $N$ individuals, but other options are equally possible. For instance, $q$ can be chosen at random from some interval, $0,q_{\mathrm{max}}] [$36], or according to some specified distribution. In the latter case, one can mimic certain network structures [31, 3, 11] in an implicit way.
@@ 5 Conclusions
For the case of a compartmental realization of this model, we found two stationary states (fixed points) for the set of differential equations of the model: the disease-free and the endemic one. They exist in their respective restricted regions of the parameter space due to the limitations for all fractions to stay positive and do not exceed $1$. The linear stability analysis indicates the stability of both fixed points within their respective regions. The expression of the basic reproductive number enables us to discuss the ways to lower the number of infected individuals in a stationary state via changing the model parameters. This can be achieved by lowering the contact rate $\beta$ and/or by the increase of the identification $\alpha$ and vaccination $\omega$ rates. However, if $\beta$ equals or exceeds the critical value $\beta_{c}$, the disease-free fixed point cannot be achieved at any combination of $\alpha$ and $\delta$. The expression for the basic reproductive number enables prediction of the most balanced approach to bring a pandemic down in economic terms, if the rates can be estimated from the real-life statistics.

### 30 Mar 2026 v2

@@ 1 Introduction and the model
The minimalistic SEIRS compartmental model, which incorporates the effects of vaccination and of a loss of immunity, is illustrated in figure 1. It contains four compartments, classified by the type of individuals they contain: susceptible to virus, $S$; unidentified infective, $E$; identified isolated infective, II; and recovered non-infective, $R$. The same letters, $S$, $E$, II, and $R$, denote respective fractions of individuals they contain.
@@
Infecting (transfer of an individual from $S$ to $E$ compartment) occurs with the rate of $\beta$. To keep the model simple, we do not introduce a separate compartment for exposed individuals [23, 48, 41, 42]. Therefore, the infecting rate, $\beta$, is affected by both social distancing, and by keeping sanitary norms, but these two factors cannot be separated in our model. Infected individual became instantly infective, as follows from the known facts of infectivity during the incubation period [66]. Infective individuals are identified (transferred from $E$ to II), via testing, with a rate of $\alpha$. Again, not to overload a model with too many processes, we consider the case of an ideally disciplined society, where the identified infective individual becomes instantly isolated, either in home or in a hospital. The difference in curing in both cases is rather minor, save for fatal and severe forms. Therefore, we do not distinguish such patients, in contrast to some other models [23, 41]. Both unidentified and isolated infective individuals are supposed to recover, acquire temporal immunity, and loose their infectivity (transferred to $R$ compartment) with the same recovery rate of $\gamma$. Vaccination provides a shortcut from the $S$ compartment straight into the $R$ one. We note, that vaccination in our model is explicit, and changed dynamically with the time evolution of the system, contrary to a static vaccination case [10], where the fraction of immune individuals is set constant. We neglect the vaccine activation period, and do not separate such factors as population coverage by vaccination and vaccine efficacy [41], since both are combined into a single vaccination rate $\omega$. In this way, all individuals residing in the $R$ compartment are assumed to be temporarily immunized. Upon loss of immunity these are transferred back into the susceptible group $S$ with the loss of immunity rate $\varphi$. These “returned” individuals may be subject to reinfection [18, 59, 27]. This may occur due to the mutation of the SARS-CoV-2 virus, e.g., into Alpha, Delta, or Omicron variant, but this can be taken into account in our model only implicitly, via the increase of the loss of immunity rate of $\varphi$. The model also neglects the birth and death rates.
@@
The rates $\beta$, $\alpha$, $\gamma$, $\varphi$ and $\omega$ can be related to the real life statistical data. If the population size is NN individuals, then the absolute number of susceptible, unidentified infective, identified isolated infective and recovered non-infective individuals are $N_{S}=SN$, $N_{E}=EN$, $N_{I}=IN$ and $N_{R}=RN$, respectively. Assuming a time unit equal to a single day, the absolute number of individuals infected per day is $\beta N_{S}(N_{E}/N)$. The coefficient $\alpha$ reflects how widely the population is covered by appropriate medical testing and defines the number of newly identified infective individuals per day, $\alpha N_{E}$. Such type of data should be available from medical institutions. Recovery rate $\gamma$, is inversely proportional to the average period of infectivity, for which much data have been systemized in reference [77]. There is some distribution of the results depending on many factors, but a reasonable average estimate of 1414 days is often cited [62, 19, 6, 69, 51]. This leads to the following estimate: $\gamma=1/14$. The loss of immunity rate, $\varphi$, inversely proportional to the time interval between (re)infections, is difficult to estimate as an unique average value [74]. Though the early studies reported the time interval for the loss of immunity of 2–3 months [4], 3–4 months [7], as well as long lasting immune memory [21], later works emphasize the role of vaccination, advanced age, the presence of comorbidities [29]. The presence of chronic diseases, namely: ischemic and inflammatory heart disease, dysrhythmias, venous thromboembolism, cerebrovascular diseases are found as an important factor as well [75]. Therefore, one may face the need to fix the time interval to a certain value, e.g., 120120 days, yielding the estimate for $\varphi=1/120$. Finally, the vaccination rate, $\omega$, can be estimated from the absolute number of individuals vaccinated per day, $\omega N_{S}$, the information available from medical institutions.
@@
One can clearly identify the effect of the presence of isolated II and recovered $R$ individuals in this model by introducing cumulative fractions of all uninfected individuals, $S^{\prime}=S+R$, and all infected ones, $E^{\prime}=E+I$, and the normalization is: $S^{\prime}+E^{\prime}=1$. The equations set (1.1)–(1.4) can be rewritten as a single equation for $E^{\prime}$
@@
which has the same form as the one for the SISSIS epidemiology model [36], when $S$ and II are substituted via $S^{\prime}$ and $E^{\prime}$, respectively, but with the reduced number of transmission acts. The latter are given by the product $(E^{\prime}-I)(S^{\prime}-R)$, where both infective and susceptible parties are lessened by II and $R$, respectively. The equation (1.6) is, obviously, not self-sufficient, as one needs to complement it by the equations providing the time evolution for the II and $R$ fractions. To do so, one may attempt either to rewrite the equations set (1.1)–(1.4) in terms of four variables $S^{\prime}$, $E^{\prime}$, II and $R$, or to use certain approximations to express $E$ and $R$ fractions via $S^{\prime}$ and $I^{\prime}$ variables, similarly as suggested in reference [37] for the SEIRS model without the immunity and vaccination considered there.
@@ 2 Fixed points and their stability
where $R_{0}$ has a meaning of the basic reproductive number. Let us note that in the limit case of $\alpha=\omega=0$ (no identification and no vaccination), the expression $R_{0}=\beta/\gamma$ for the SIS model is retrieved (in this case the $E$ fraction of the current model serves as the II fraction in the SIS model). Taking into account the assumption (2.6), the expression (2.16) indicates the way of bringing the basic reproductive number $R_{0}$ down by means of both decrease of the transmission rate $\beta$ and by an increase of the identification $\alpha$ and vaccination $\omega$ rates. One can rewrite the condition $R_{0}\leqslant 1$ for the DF state in any of the following three forms
@@
at the time instance $t$ depend on all variables, $S$, $E$, II and $R$ at the same instance $t$. The source of infection is the $E$ fraction, because in the current model, depicted in figure 1, we consider an ideal case of instant isolation of identified infective individuals into a group II. On the other hand, part of the latter are isolated in the hospitals and at home, therefore II reflects the load put on the health system. Therefore, we concentrate on the time evolutions, $E(t)$ and $I(t)$, of these two fractions. These are examined at various initial values of $E_{0}$ and at various $\beta$, $\alpha$ and $\omega$ rates.
@@ 4 Cellular automaton simulations
To take into account the spatial arrangement of individuals within the SEIRS model as defined in figure 1, we consider their simplest 2D arrangement: a simple square lattice with a neighbourhood size for each individual, given by the number $q$ of its neighbours. The setup is similar to that being discussed for the case of a simpler SIS model in reference [36]. In particular, the vertex set of a graph is $\mathbb{Z}^{2}$, and each individual is characterized by a neighbourhood radius $R_{q}>0$, such that the neighbourhood of a given $k\in\mathbb{Z}^{2}$ is defined as: $k^{\prime}\sim k$ whenever $|k^{\prime}-k|\leqslant R_{q}$. The choice of $R_{q}$ is made to obtain the neighbourhood size equal to the previously chosen neighbourhood size $q$. Each $k$ is associated with an individual, that is characterized by its state $s_{k}$, which takes any value from a set of $\{s,e,i,r\}$, the letters correspond to the susceptible, unidentified infected, identified infected and recovered states, respectively. The evolution of the system of NN such individuals is driven by the algorithm, based on transition rules between the states of each individual in the SEIRS model, depicted schematically in figure 1. The principal difference with the compartmental realization of this model is that the transmission occurs locally, within the sphere of the radius $R_{q}$ around each infected individual.
We consider a square lattice of $N=700^{2}=490\,000$ individuals which brings the model to a realistic size of a medium-sized town. The periodic boundary conditions are applied along both Cartesian axes. The set of neighbourhood sizes ranges from $q=N$ (the limit of the compartmental model) down to $q=4$ (von Neumann neighbourhood). In this way the effect of quarantine is considered explicitly, by reducing the neighbourhood size $q$, and can be partly separated from the probability to contract a disease. At each time instance, we made NN random choices of individuals, and their states are updated according to their current state and the states of their neighbours. This levels up the time discretization with the $\Delta t$ of the numeric integration (3.1). We employed the asynchronous realization of the cellular automaton algorithm, when the change of $s_{k}$ occurs immediately, and affects the rest of NN updating attempts performed at the same time instance.
We focus here on the time evolution of the fraction of identified infected individuals, $I(t)$, as it characterizes the load put on the medical system of an imaginary town. The strength of quarantine measures is defined by restrictions imposed on social contacts defined via the neighbourhood size $q$. It is fixed throughout the system of NN individuals, but other options are equally possible. For instance, $q$ can be chosen at random from some interval, $0,q_{\mathrm{max}}] [$36], or according to some specified distribution. In the latter case, one can mimic certain network structures [31, 3, 11] in an implicit way.
@@ 5 Conclusions
For the case of a compartmental realization of this model, we found two stationary states (fixed points) for the set of differential equations of the model: the disease-free and the endemic one. They exist in their respective restricted regions of the parameter space due to the limitations for all fractions to stay positive and do not exceed 11. The linear stability analysis indicates the stability of both fixed points within their respective regions. The expression of the basic reproductive number enables us to discuss the ways to lower the number of infected individuals in a stationary state via changing the model parameters. This can be achieved by lowering the contact rate $\beta$ and/or by the increase of the identification $\alpha$ and vaccination $\omega$ rates. However, if $\beta$ equals or exceeds the critical value $\beta_{c}$, the disease-free fixed point cannot be achieved at any combination of $\alpha$ and $\delta$. The expression for the basic reproductive number enables prediction of the most balanced approach to bring a pandemic down in economic terms, if the rates can be estimated from the real-life statistics.